An air bag needs 60 dm3 of gas to inflate it at 25°C. Calculate: 1.1 the number of moles of N2 gas at 298 K required to inflate the air bag to a pressure of 1 atm. (7) 1.2 the number of moles of NaN3 required to produce this amount of gas (Hint: NaN3 reacts to produce Na(s) and N2(g). (5) 1.3 the mass of NaN3 required.
1
Expert's answer
2020-12-09T06:34:05-0500

2Na3N = 6Na + N2

PV = nRT

n = PV/RT

= 1 ×60 / 0.082 × 298

= 2.46moles

2. Mole of Na3N = 2.46/2 = 1.23 moles

3. Mass of Na3N = 1.23 × (83) = 102.09g